### v5

# DRCT: Saving Image Super-Resolution away from Information Bottleneck

Abstract
In recent years, Vision Transformer-based approaches for low-level vision tasks have achieved widespread success. Unlike CNN-based models, Transformers are more adept at capturing long-range dependencies, enabling the reconstruction of images utilizing non-local information. In the domain of super-resolution, Swin-transformer-based models have become mainstream due to their capability of global spatial information modeling and their shifting-window attention mechanism that facilitates the interchange of information between different windows. Many researchers have enhanced model performance by expanding the receptive fields or designing meticulous networks, yielding commendable results. However, we observed that it is a general phenomenon for the feature map intensity to be abruptly suppressed to small values towards the network’s end. This implies an information bottleneck and a diminishment of spatial information, implicitly limiting the model’s potential. To address this, we propose the Dense-residual-connected Transformer (DRCT), aimed at mitigating the loss of spatial information and stabilizing the information flow through dense-residual connections between layers, thereby unleashing the model’s potential and saving the model away from information bottleneck. Experiment results indicate that our approach surpasses state-of-the-art methods on benchmark datasets and performs commendably at the NTIRE-2024 Image Super-Resolution (x4) Challenge. Our source code is available at https://github.com/ming053l/DRCT
1 Introduction
Figure 1: The feature map intensity on various benchmark datasets. We observed that feature map intensities decrease sharply at the end of SISR network, indicating potential information loss. In this paper, we propose DRCT to address this issue by enhancing receptive fields and adding dense-connections within residual blocks to mitigate information bottlenecks, thereby improving performance with a simpler model design.
The task of Single Image Super-Resolution (SISR) is aimed at reconstructing a high-quality image from its low-resolution version. This quest for effective and skilled super-resolution algorithms has become a focal point of research within the field of computer vision, owing to its wide range of applications.
Following the foundational studies, CNN-based strategies [8, 35, 20, 19, 40, 39] have predominantly governed the super-resolution domain for an extended period. These strategies largely leverage techniques such as residual learning [61, 52, 46, 12, 36], or recursive learning [21, 18] for developing network architectures, significantly propelling the progress of super-resolution models forward.
CNN-based networks have achieved notable success in terms of performance. However, the inductive bias of CNN limits SISR models capture long-range dependencies. Their inherent limitations stem from the parameter-dependent scaling of the receptive field and the kernel size of convolution operator within different layers, which may neglect non-local spatial information within images.
To overcome the limitations associated with CNN-based networks, researchers have introduced Transformer-based SISR networks that leverage the capability to model long-range dependencies, thereby enhancing SISR performance. Notable examples include IPT [16] and EDT [33], which utilize pre-training on large-scale dataset like ImageNet [14] to fully leverage the capabilities of Vision Transformer [9] for achieving ideal SISR results. Afterwards, SwinIR [34] incorporates Swin-Transformer [26] into SISR, marked a significant advancement in SISR performance.
This approach significantly enhances capabilities beyond those of traditional CNN-based models across various benchmarks. Following SwinIR’s success, several works [63, 6, 4, 64, 34, 32, 58, 59] have built upon its framework. These subsequent studies leverage Transformers to innovate diverse network architectures specifically for super-resolution tasks, showcasing the evolving landscape of SISR technology through the exploration of new architectural innovations and techniques.
While using Transformer-based SISR model for inference across various datasets, we observed a common phenomenon: the intensity distribution of the feature maps undergoes more substantial changes as the network depth increases. This indicates the spatial information and attention intensity learned by the model. However, there’s often sharp decrease towards the end of the network (refer to Figure 1) , shrinking to a smaller range. This phenomenon suggests that such abrupt changes might be accompanied by a loss of spatial information, indicating the presence of an information bottleneck.
Inspired by a series of works by Wang et al., such as the YOLO-family [50, 47], CSPNet [48], and ELAN [49], we consider that network architectures based on SwinIR, despite significantly enlarging the receptive fields through shift-window attention mechanism to address the small receptive fields in CNNs, are prone to gradient bottlenecks due to the loss of spatial information as network depth increases. This implicitly constrains the model’s performance and potential.
To address the issue of spatial information loss due to an increased number of network layers, we introduce the Dense-residual-connected Transformer (DRCT), designed to stabilize the forward-propagation process and prevent information bottlenecks. This is achieved by the proposed Swin-Dense-Residual-Connected Block (SDRCB), which incorporates Swin Transformer Layers and transition layers into each Residual Dense Group (RDG). Consequently, this approach enhances the receptive field with fewer parameters and a simplified model architecture, thereby resulting in improved performance. The main contributions of this paper are summarised as follows:
  • We observed that as the network depth increases, the intensity of the feature map will gradually increase, then abruptly drop to a smaller range. This severe oscillation may be accompanied by information loss.
  • We propose DRCT by adding dense-connection within residual groups to stabilize the information flow for deep feature extraction during propagation, thereby saving the SISR model away from the information bottleneck.
  • By integrating dense connections into the Swin-Transformer-based SISR model, the proposed DRCT achieves state-of-the-art performance while maintaining efficiency. This approach showcases its potential and raises the upper-bound of the SISR task.
2 Related works
2.1 Vision Transformer-based Super-Resolution
IPT [16], a versatile model utilizing the Transformer encoder-decoder architecture, has shown efficacy in several low-level vision tasks. SwinIR [34], building on the Swin Transformer [26] encoder, employs self-attention within local windows during feature extraction for larger receptive fields and greater performance, compared to traditional CNN-based approaches. UFormer [53] introduces an innovative local-enhancement window Transformer block, utilizing a learnable multi-scale restoration modulator within the decoder to enhance the model’s ability to detect both local and global patterns. ART [64] incorporates an attention retractable module to expand its receptive field, thereby enhancing SISR performance. CAT [5] leverages rectangle-window self-attention for feature aggregation, achieving a broader receptive field. HAT [4] integrates channel attention mechanism [51] with overlapping cross-attention module, activating more pixels to reconstruct better SISR results, thereby setting new benchmarks in the field.
Figure 2: The feature map visualization displays, from top to bottom, SwinIR [34], HAT [4], and the proposed DRCT, with positions further to the right representing deeper layers within the network. For both SwinIR and HAT, the intensity of the feature maps is significant in the shallower layers but diminishes towards the network’s end. We consider this phenomenon implies the loss of spatial information, leading to the limitation and information bottleneck with SISR tasks. As for the proposed DRCT, the learned feature maps are gradually and stably enhanced without obvious oscillations. It represents the stability of the information flow during forward propagation, thereby yielding higher intensity in the final layer’s output. (zoom in to better observe the color-bar besides feature maps.)
2.2 Auxiliary Supervision and Feature Fusion
Auxiliary Supervision. Deep supervision is a commonly used auxiliary supervision method [31, 13] that involves training by adding prediction layers at the intermediate levels of the model [47, 48, 49]. This approach is particularly prevalent in architectures based on Transformers that incorporate multi-layer decoders. Another popular auxiliary supervision technique involves guiding the feature maps produced by the intermediate layers with relevant metadata to ensure they possess attributes beneficial to the target task [11, 28, 52, 61, 30]. Choosing the appropriate auxiliary supervision mechanism can accelerate the model’s convergence speed, while also enhancing its efficiency and performance.
Feature Fusion. Many studies have explored the integration of features across varying dimensions or multi-level features, such as FPN [22], to obtain richer representations for different tasks [36, 55]. In CNNs, attention mechanisms have been applied to both spatial and channel dimensions to improve feature representation; examples of which include RTCS [10] and SwinFusion [37]. In ViT [9], spatial self-attention is used to model the long-range dependencies between pixels. Additionally, some researchers have investigated the incorporation of channel attention within Transformers [62, 3] to effectively amalgamate spatial and channel information, thereby improving model performance.
3 Problem Statement
3.1 Information Bottleneck Principle
According to the information bottleneck principle [45], the given data $X$ may cause information loss when going through consecutive layers. It may lead to gradient vanish when back-propagation for fitting network parameters and predicting $Y$, as shown in the equation below:
$I(X,X)\geq I(Y,X)\geq I(Y,f_{\theta}(X))\geq I(X,g_{\phi}(f_{\theta}(X))),$ (1)
where $I$ indicates mutual information, $f$ and $g$ are transformation functions, and $\theta$ and $\phi$ are parameters of $f$ and $g$, respectively.
In deep neural networks, $f_{\theta}(\cdot)$ and $g_{\phi}(\cdot)$ respectively represent the two consecutive layers in neural network. From equation (1), we consider that as the number of network layer becomes deeper, the information flow will be more likely to be lost. In term of SISR tasks, the general goal is to find the mapping function $F$ with optimized function parameters $\theta$ to maximize the mutual information between HR and SR image.
$F(\mathbf{I}_{LR};\theta)=\mathbf{I}_{SR};\\ \max_{\theta}I(\mathbf{I}_{HR};F(\mathbf{I}_{LR};\theta))$ (2)
3.2 Spatial Information Vanish in Super-resolution
Generally speaking, SISR methods [4, 5, 34, 58, 64, 6, 63, 32] can generally divided into three parts: (1) shallow feature extraction, (2) deep feature extraction, (3) image reconstruction. Among these methods, there is almost no difference between shallow feature extraction and image reconstruction. The former is composed of simple convolution layers, while the latter consists of convolution layers and upsampling layers. However, deep feature extraction differs significantly. Yet, their commonality lies in being composed of various residual blocks, which can be simply defined as:
$X^{l+1}=X^{l}+f_{\theta}^{l+1}(X^{l}),$ (3)
where $X$ indicates inputs, $f$ is a consecutive layers for $l$’th residual group , and $\theta$ represents the parameters of $f^{l}$.
Especially for SISR task, two methods of stabilizing information flow or training process are introduced:
Residual connection to learn local feature. Adopting residual learning allows the model to only update the differences between layers, rather than output the total information from a previous layer directly [12]. This reduces the difficulty of model training and prevents gradient vanishing locally [61]. However, according to our observations, while this design effectively transmits spatial information between different residual blocks, there may still be information loss.
Because the information within a residual block may not necessarily maintain spatial information, this ultimately leads to non-smoothness in terms of feature map intensity (refer to Fig. 2), causing an information bottleneck at the deepest layers during forward propagation. This makes it easy for spatial information to be lost as the gradient flow reaches the deeper layers of the network, resulting in reduced data efficiency or the need for more complex network designs to achieve better performance.
Dense connection to stabilize information flow. Incorporating dense connections into the Swin-Transformer based SISR model offers two significant advantages. Firstly, global auxiliary supervision. It effectively fuses the spatial information across different residual groups [52, 61], preserving high-frequency features throughout the deep feature extraction. Secondly, saving SISR model away from information bottleneck. By leveraging the integration of spatial information, the model ensures a smooth transmission of spatial information [46], thereby mitigating the information loss and enhancing the receptive field.
4 Motivation
Dense-Residual Group auxiliary supervision. Motivated by RRDB-Net [52], Wang et al. suggested that incorporating dense-residual connections can aggregate multi-level spatial information and stabilize the training process [35, 41]. We consider that it is possible to stabilize the information flow within each residual-groups during propagation, thereby saving SISR model away from the information bottleneck.
Dense connection with Shifting-window mechanism. Recent studies on SwinIR-based methods have concentrated on enlarging the receptive field [4, 5, 64, 6] by sophisticated WSA or enhancing the network’s capability to extract features [32, 53] for high-quality SR images. By adding dense-connections [15] within Swin-Transformer-based blocks [34, 26] in the SISR network for deep feature extraction, the proposed DRCT’s receptive field is enhanced while capturing long-range dependencies. Consequently, this approach allows for achieving outstanding performance with simpler model architectures [46], or even using shallower SISR networks.
5 Methodology
5.1 Network Architecture
Figure 3: The overall architecture of the proposed Dense-residual-connected Transformer (DRCT) and the structure of Residual-Dense Group (RDG). Each RDG contains five consecutive Swin-Dense-Residual-Connected Blocks (SDRCBs). By integrating dense-connection [15] into SwinIR [34], the efficiency can be improved for Saving Image Super-resolution away from Information Bottleneck.
As shown in Figure 3, DRCT comprises three distinct components: shallow feature extraction, deep feature extraction, and image reconstruction module, respectively.
Shallow and deep feature extraction. Given a low-resolution (LR) input $\mathbf{I}_{LR}\in\mathbb{R}^{H\times W\times C_{in}}$ ($H$, $W$ and $C_{in}$ are the image height, width and input channel number, respectively), we use a $3\times 3$ convolution layer Conv()($\cdot$)( ⋅ ) [54] to extract shallow feature $\mathbf{F}_{\text{0}}\in\mathbb{R}^{H\times W\times C}$ as
$\mathbf{F}_{\text{0}}=\text{Conv}(\mathbf{I}_{LQ}),$ (4)
Then, we extract deep feature which contains high-frequency spatial information $\mathbf{F}_{DF}\in\mathbb{R}^{H\times W\times C}$ from $\mathbf{F}_{\text{0}}$ and it can be defined as
$\mathbf{F}_{DF}=H_{DF}(\mathbf{F}_{0}),$ (5)
where $H_{DF}(\cdot)$ is the deep feature extraction module and it contains $K$ Residual Dense Group (RDG) and single convolution layer $\text{Conv}(\cdot)$ for feature transition. More specifically, intermediate features $\mathbf{F}_{1},\mathbf{F}_{2},\ldots,\mathbf{F}_{K}$ and the output deep feature $\mathbf{F}_{DF}$ are extracted block by block as
$\mathbf{F}_{i}=\text{RDG}_{i}(\mathbf{F}_{i-1}),\quad i=1,2,\ldots,K,$ (6)
$\mathbf{F}_{DF}=\text{Conv}(\mathbf{F}_{K}),$ (7)
Image reconstruction. We reconstruct the SR image $\mathbf{I}_{SR}$ $\in\mathbb{R}^{H\times W\times C_{in}}$ by aggregating shallow and deep features, it can be defined as:
$\mathbf{I}_{SR}=H_{\text{rec}}(\mathbf{F}_{0}+\mathbf{F}_{DF}),$ (8)
where $H_{\text{rec}}(\cdot)$ is the function of the reconstruction for fusing high-frequency deep feature $\mathbf{F}_{DF}$ and low-frequency feature $\mathbf{F}_{0}$ together to obtain SR result.
5.2 Deep Feature Extraction
Residual Dense Group. In developing of RDG, we take cues from RRDB-Net [52] and RDN [61], employing a residual-dense block (RDB) as the foundational unit for SISR. The reuse of feature maps emerges as the enhanced receptive field in the RDG’s feed-forward mechanism. To expound further, RDG with several SDRCB enhances the capability to integrate information across different scales, thus allowing for a more comprehensive feature extraction. RDG facilitates the information flow within residual group, capturing the local features and spatial information group by group.
Swin-Dense-Residual-Connected Block. In purpose of capturing the long-range dependency, we utilize the shifting window self-attention mechanism of Swin-Transformer Layer (STL) [26, 34] for obatining adaptive receptive fields, complementing RRDB-Net by focusing on multi-level spatial information. This synergy leverages STL to dynamically adjust the focus of the model based on the global content of the input, allowing for a more targeted and efficient extraction of features. This mechanism ensures that even as the depth of the network increases, global details are preserved, effectively enlarging and enhancing the receptive field without compromising. By integrating STLs with dense-residual connections, the architecture benefits from both a vast receptive field and the capability to hone in on the most relevant information, thereby enhancing the model’s performance in SISR tasks requiring detailed and context-aware processing. For the input feature maps $\mathbf{Z}$ within RDG, the SDRCB can be defined as:
$\mathbf{Z}_{j}=H_{\text{trans}}(\text{STL}([\mathbf{Z},...,\mathbf{Z}_{j-1}])),j=1,2,3,4,5,$ (9)
$\text{SDRCB}(\mathbf{Z})=\alpha\cdot\mathbf{Z}_{5}+\mathbf{Z},$ (10)
where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation of multi-level feature maps produced by the previous layers. $H_{\text{trans}}(\cdot)$ refers to the convolution layer with a LeakyReLU activate function for feature transition. The negative slope of LeakyReLU is set to $0.2$. $\text{Conv}_{1}$ is the $1\times 1$ convolution layer, which is used to adaptively fuse a range of features with different levels [42]. $\alpha$ represents residual scaling factor, which is set to $0.2$ for stabilizing the training process [52].
5.3 Same-task Progressive Training Strategy
In recent years, Progressive Training Strategy (PTS) [17, 25] has gained increased attention and can be seen as a method of fine-tuning. Compared to conventional training methods, PTS tends to converge model parameters to more desirable local minima. HAT [4] introduces the Same-task Pre-training, which aims to train the model on a large dataset like ImageNet [14] before fine-tuning it on a specific dataset, leading to improved SISR results. Lei et al. [57] proposed initially training a SISR network with L1-loss and then using L2-loss to eliminate artifacts, achieving better results on the PSNR metric. This has been widely adopted [64]. We proposed a Same-task Progressive Training Strategy (SPTS). At first, we pre-trained DRCT on ImageNet to initialize model parameters and then fine-tuned on specific datasets with L1 loss,
$\ell_{L1}=\left\|I_{HR}-I_{SR}\right\|_{1},$ (11)
and finally use L2 loss to eliminate singular pixels and artifacts, therefore further archiving greater performance on PSNR metric.
$\ell_{L2}=\left\|I_{HR}-I_{SR}\right\|_{2}$ (12)
6 Experiment Results
Method Scale Training Set5 [2] Set14 [56] BSD100 [38] Urban100 [29] Manga109 [24]
Dataset PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
EDSR [35] ×$\times 2$ DIV2K 38.11 0.9602 33.92 0.9195 32.32 0.9013 32.93 0.9351 39.10 0.9773
RCAN [60] ×$\times 2$ DIV2K 38.27 0.9614 34.12 0.9216 32.41 0.9027 33.34 0.9384 39.44 0.9786
SAN [20] ×$\times 2$ DIV2K 38.31 0.9620 34.07 0.9213 32.42 0.9028 33.10 0.9370 39.32 0.9792
IGNN [19] ×$\times 2$ DIV2K 38.24 0.9613 34.07 0.9217 32.41 0.9025 33.23 0.9383 39.35 0.9786
HAN [40] ×$\times 2$ DIV2K 38.27 0.9614 34.16 0.9217 32.41 0.9027 33.35 0.9385 39.46 0.9785
NLSN [39] ×$\times 2$ DIV2K 38.34 0.9618 34.08 0.9231 32.43 0.9027 33.42 0.9394 39.59 0.9789
SwinIR [34] ×$\times 2$ DF2K 38.42 0.9623 34.46 0.9250 32.53 0.9041 33.81 0.9427 39.92 0.9797
CAT-A [5] ×$\times 2$ DF2K 38.51 0.9626 34.78 0.9265 32.59 0.9047 34.26 0.9440 40.10 0.9805
HAT [4] ×$\times 2$ DF2K 38.63 0.9630 34.86 0.9274 32.62 0.9053 34.45 0.9466 40.26 0.9809
DAT [6] ×$\times 2$ DF2K 38.58 0.9629 34.81 0.9272 32.61 0.9051 34.37 0.9458 40.33 0.9807
DRCT (Ours) ×$\times 2$ DF2K 38.62 0.9628 34.84 0.9272 32.62 0.9051 34.44 0.9464 40.31 0.9804
[…]
EDT† [33] ×$\times 2$ DF2K 38.63 0.9632 34.80 0.9273 32.62 0.9052 34.27 0.9456 40.37 0.9811
[…]
EDSR [35] ×$\times 3$ DIV2K 34.65 0.9280 30.52 0.8462 29.25 0.8093 28.80 0.8653 34.17 0.9476
RCAN [60] ×$\times 3$ DIV2K 34.74 0.9299 30.65 0.8482 29.32 0.8111 29.09 0.8702 34.44 0.9499
SAN [20] ×$\times 3$ DIV2K 34.75 0.9300 30.59 0.8476 29.33 0.8112 28.93 0.8671 34.30 0.9494
IGNN [19] ×$\times 3$ DIV2K 34.72 0.9298 30.66 0.8484 29.31 0.8105 29.03 0.8696 34.39 0.9496
HAN [40] ×$\times 3$ DIV2K 34.75 0.9299 30.67 0.8483 29.32 0.8110 29.10 0.8705 34.48 0.9500
NLSN [39] ×$\times 3$ DIV2K 34.85 0.9306 30.70 0.8485 29.34 0.8117 29.25 0.8726 34.57 0.9508
SwinIR [34] ×$\times 3$ DF2K 34.97 0.9318 30.93 0.8534 29.46 0.8145 29.75 0.8826 35.12 0.9537
CAT-A [5] ×$\times 3$ DF2K 35.06 0.9326 31.04 0.8538 29.52 0.8160 30.12 0.8862 35.38 0.9546
HAT [4] ×$\times 3$ DF2K 35.07 0.9329 31.08 0.8555 29.54 0.8167 30.23 0.8896 35.53 0.9552
DAT [6] ×$\times 3$ DF2K 35.16 0.9331 31.11 0.8550 29.55 0.8169 30.18 0.8886 35.59 0.9554
DRCT (Ours) ×$\times 3$ DF2K 35.18 0.9338 31.24 0.8569 29.68 0.8182 30.34 0.8910 35.76 0.9575
[…]
EDT† [33] ×$\times 3$ DF2K 35.13 0.9328 31.09 0.8553 29.53 0.8165 30.07 0.8863 35.47 0.9550
[…]
EDSR [35] ×$\times 4$ DIV2K 32.46 0.8968 28.80 0.7876 27.71 0.7420 26.64 0.8033 31.02 0.9148
RCAN [60] ×$\times 4$ DIV2K 32.63 0.9002 28.87 0.7889 27.77 0.7436 26.82 0.8087 31.22 0.9173
SAN [20] ×$\times 4$ DIV2K 32.64 0.9003 28.92 0.7888 27.78 0.7436 26.79 0.8068 31.18 0.9169
IGNN [19] ×$\times 4$ DIV2K 32.57 0.8998 28.85 0.7891 27.77 0.7434 26.84 0.8090 31.28 0.9182
HAN [40] ×$\times 4$ DIV2K 32.64 0.9002 28.90 0.7890 27.80 0.7442 26.85 0.8094 31.42 0.9177
NLSN [39] ×$\times 4$ DIV2K 32.59 0.9000 28.87 0.7891 27.78 0.7444 26.96 0.8109 31.27 0.9184
SwinIR [34] ×$\times 4$ DF2K 32.92 0.9044 29.09 0.7950 27.92 0.7489 27.45 0.8254 32.03 0.9260
CAT-A [5] ×$\times 4$ DF2K 33.08 0.9052 29.18 0.7960 27.99 0.7510 27.89 0.8339 32.39 0.9285
HAT [4] ×$\times 4$ DF2K 33.04 0.9056 29.23 0.7973 28.00 0.7517 27.97 0.8368 32.48 0.9292
DAT [6] ×$\times 4$ DF2K 33.08 0.9055 29.23 0.7973 28.00 0.7515 27.87 0.8343 32.51 0.9291
DRCT (Ours) ×$\times 4$ DF2K 33.11 0.9064 29.35 0.7984 28.18 0.7532 28.06 0.8378 32.59 0.9304
[…]
EDT† [33] ×$\times 4$ DF2K 32.82 0.9031 29.09 0.7939 27.91 0.7483 27.46 0.8246 32.05 0.9254
[…]
Table 1: Quantitative comparison with the several peer-methods on benchmark datasets. $\dagger$” indicates that methods adopt pre-training strategy [4] on ImageNet. $\ddagger$” represents that methods use same-task progressive-training strategy. The top three results are marked in red, blue, and orange , respectively.
Figure 4: Visual comparison on × 4 SISR. The patches for comparison are marked with red boxes in the original images. The higher the PSNR/SSIM metrics, the better the performance..
6.1 Dataset
Our DRCT model is trained on DF2K, a substantial aggregated dataset that includes DIV2K [1] and Flickr2K [44]. DIV2K provides 800 images for training, while Flickr2K contributes 2650 images. For the training input, we generate LR versions of these images by applying a bicubic downsampling method with scaling factors of $2$, $3$, and $4$, respectively. To assess the effectiveness of our model, we conduct performance evaluations using well-known SISR benchmark datasets such as Set5 [2], Set14 [56], BSD100 [38], Urban100 [29], and Manga109 [24].
6.2 Implementation Details
The training process can be structured into three phases, as Section 4-3 illustrates. (1) pre-trained on ImageNet [14], (2) optimize the model on the given dataset, (3) L2-loss for PSNR enhancement. Throughout the training process, we use the Adam optimizer with $\beta_{1}$ = $0.9$, and $\beta_{2}$ = $0.999$ and train for $800k$ iterations in the first and second stages. The learning rate is set to $42e-$ 4, and the multi-step learning scheduler is also used. The learning rate is halved at the $300k$, $500k$, $650k$, $700k$, $750k$ iterations respectively. Weight decay is not applied, and the batch size is set to $32$. In the architecture of DRCT, the configuration of depth and width is maintained identically to that of HAT [4]. To elaborate, both the number of RDG and SDRCB units are established at 6, and the channel number of intermediate feature maps is designated as 180. The attention head number and window size are set to 6 and 16 for window-based multi-head self-attention (W-MSA). In terms of data preparation, HR patches with dimensions of $256$$\times$ $256$ pixels were extracted from the HR images. To improve the generalizability, we apply random horizontal flips and rotation augmentation.
6.3 Quantitative Results
For the evaluation, we use full RGB channels and ignore the $(2\times\text{scale})$ pixels from the border. PSNR and SSIM metrics are used to evaluation criteria. Table 1 presents the quantitative comparison of our approach and the state-of-the-art methods, including EDSR [35], RCAN [60], SAN [20], IGN [19], HAN [40], NLSN [39], SwinIR [34], CAT-A [5], DAT [6], as well as approaches using ImageNet pre-training, such as IPT [16], EDT [33] and HAT [4]. We can see that our method outperforms the other methods significantly on all benchmark datasets. In addition, the DRCT-L can bring further improvement and greatly expand the performance upper-bound on SISR tasks. Even with fewer model parameters and computational requirements, DRCT is also significantly greater than the state-of-the-art methods.
6.4 Visual Comparison
The visual comparisons displayed in Figure 4. For the selected images from Urban100 [29], DRCT is effective in restoring structures, whereas other methods suffer from notably blurry effects. The visual results demonstrate the superiority of our approach.
Along with providing visualizations for the LAM [27], we compute the Diffusion Index (DI), which is the attribution-based analysis. The DI reflects the range of involved pixels. A higher DI refers to a wider range of attention. In scenarios where DRCT used fewer parameters (which will be discussed in the next subsection), it achieves a higher DI. This outcome suggests that, after enhancing the receptive field through SDRCB, the model can leverage a long-range dependency and non-local information for SISR without the need for intricately designed W-MSA.
Figure 5: The LAM [27] visualization. DRCT improves performance by enhancing the receptive field to mitigate the issue of spatial information loss in deeper layers of the network.
Figure 6: The model complexity comparison between SwinIR, HAT, and proposed DRCT evaluated on Urban100 [29] dataset.
6.5 Model Complexity
To demonstrate the potential of our proposed DRCT, we conducted further analysis on model complexity and performance.
Model efficiency. In Table 2, the proposed DRCT clearly requires fewer computational resources compared to HAT in terms of parameter size, multiply-add operations, memory requirements, and FLOPs. Specifically, when scaling up the model sizes of DRCT and HAT, DRCT-L surpasses HAT-L in all metrics.
Model performance. From Figure 6, we can observe that the performance curves of the HAT and SwinIR models are approaching horizontal lines, suggesting that the performance is nearing a bottleneck and its upper-bound, even if scaling up the model parameters.
This demonstrates that the design of DRCT, which incorporates dense-connections in the residual groups within a Swin-transformer-based model to stabilize the information flow, achieves convincing results with a reduced computational burden.
##\##Params. ##\##Multi-Adds. Forward or Backward pass FLOPs
HAT [4] 20.77M 11.22G 2053.42M 42.18G
DRCT 14.13M 5.92G 1857.55M 7.92G
HAT-L [4] 40.846M 76.69G 5165.39M 79.60G
DRCT-L 27.580M 9.20G 4278.19M 11.07G
Table 2: Model complexity analysis for (×4) SISR on Urban100.
6.6 NTIRE Image Super-Resolution (x4) Challenge
Validation phase Testing phase
PSNR 31.1820 31.1776
SSIM 0.8494 0.8620
Table 3: NTIRE 2024 Challenge Results with x4 SR in terms of PSNR and SSIM on validation phase and testing phase.
The dataset for the NTIRE 2024 Image Super-Resolution (x4) Challenge [7] comprises three collections: DIV2K [1], Flickr2K [44], and LSDIR [23]. Specifically, the DIV2K dataset provides 800 pairs of HR and LR images for training. The LR images are obtained from the HR images after bicubic downsampling with specific scaling factor. For validation, it offers 100 LR images for the purpose of creating SR images, with the HR versions to be made available at the challenge’s final stage. Additionally, the test dataset includes 100 varied LR images. The self-ensemble strategy is used for testing-time augmentation (TTA) [43]. Our TTA methods include random rotation, and horizontal and vertical flipping. We also conducted a model ensemble strategy for fusing different reconstructed results by HAT [4] and the proposed DRCT to eliminate the annoying artifacts and improve final SR quality. Our SISR model was entered into both the validation and testing phases of this challenge, with the detailed in Table 3.
7 Conclusion
In this paper, we introduce the phenomenon of information bottlenecks observed in SISR models, where spatial information is lost as network depth increases during forward propagation. This may lead to information loss when limiting the upper bound of model performance for the SISR task, which requires detailed spatial information and context-aware processing.
To address these issues, we present a novel Swin-transformer-based model, Dense-residual-connected Transformer (DRCT). The design philosophy behind DRCT centers on stabilizing the information flow and enhancing the receptive fields by incorporating dense-connections within residual blocks, combining the shift-window attention mechanism to adaptively capture global information.
As a result, the model can better focus on global spatial information and surpass existing state-of-the-art methods without the need for designing sophisticated window attention mechanisms or increasing model parameters. The experiment results have demonstrated the efficacy of the proposed DRCT, indicating its effectiveness and the potential for future work related to SISR tasks.